OpenStudy (anonymous):

solve for w: 2w^2+10=0

OpenStudy (anonymous):

2w^2 = -10 w^2 = -5 w = +/- root(-5) w = +/- iroot5 i.e. \[w = \pm i \sqrt{5}\]

OpenStudy (anonymous):

thanks